lat A be avector parallel to line of intersection of two planes P1 and P2 through origin. P1 is parallel to the vectors 2i+3k and 4j-3k and P2 is parallel to j-k and 3i+3j, then the angle between vector A and 2i+j-2k. ANS; 45 and 135 degree

lat  A be avector parallel to line of intersection of two planes P1 and P2 through origin. P1 is parallel to the vectors 2i+3k and 4j-3k and P2 is parallel to j-k and 3i+3j, then the angle between vector A and 2i+j-2k.


 


 


 


ANS; 45 and 135 degree

Grade:10

1 Answers

vikas askiitian expert
13 years ago

P1 is parallel to vector 2i+3j & 4j-3k

normal of P1 = (2i+3j)*(4j-3k)

                   =-9i+6j+8k (let it be B)

normal of plane P2 = (j-k)*(3i+3j)

                           =-3i+3j+3k  (let it be C)

line of intersection of planes will be B*C

     A = B*C

        =(-9i+6j+8k)*(-3i+3j+3k)

        = -6i-3j-9k

angle bw A & 2i+j-2k is

  cos@ = 1/(126)1/2      or 

      @ = cos-1 (1/3sqrt14)